A 3.0 g bullet traveling at 300 m/s hits and embeds itself in a 1.0 kg wooden block resting on a frictionless surface. What is the final speed of the block and bullet? *




1
Expert's answer
2021-04-26T17:09:46-0400
"mV=(m+M)v\\(0.003)(300)=(1+0.003)v\\v=0.90\frac{m}{s}"